The width of a rectangle is 1cm more than half the length. When both the length and the width are increased by 1cm, the area increases by 20cm^2. Find the original dimensions of the rectangle.
Answer by stanbon(75887)

The width of a rectangle is 1cm more than half the length. When both the length and the width are increased by 1cm, the area increases by 20cm^2. Find the original dimensions of the rectangle.
----
Original DATA:
length = L
width = (1/2)L+1
Area = (1/2)L^2 + L
------------------------
New DATA:
Length = L+1
width = (1/2)L+2
Area = (L+1)[(1/2)L+2]
----
Equation:
New area - Old area = 20 cm^2
(L+1)((1/2)L+2) - [(1/2)L^2+L] = 20
-----
(1/2)L^2 + (5/2)L +2 - (1/2)L^2 - L = 20
-------
(7/2)L = 18
--
L = 36/7 cm (original Length)
Original width = (1/2)L + 1 = (18/7)+(7/7) = 25/7 (original width)
=================
Cheers,
Stan H.
=================